# Sortblock: Similarity-Aware Feature Reuse for Diffusion Model

### v1

@@ Abstract
Diffusion Transformers (DiTs) have demonstrated remarkable generative capabilities, particularly benefiting from Transformer architectures that enhance visual and artistic fidelity. However, their inherently sequential denoising process results in high inference latency, limiting their deployment in real-time scenarios. Existing training-free acceleration approaches typically reuse intermediate features at fixed timesteps or layers, overlooking the evolving semantic focus across denoising stages and Transformer blocks.To address this, we propose Sortblock, a training-free inference acceleration framework that dynamically caches block-wise features based on their similarity across adjacent timesteps. By ranking the evolution of residuals, Sortblock adaptively determines a recomputation ratio, selectively skipping redundant computations while preserving generation quality. Furthermore, we incorporate a lightweight linear prediction mechanism to reduce accumulated errors in skipped blocks.Extensive experiments across various tasks and DiT architectures demonstrate that Sortblock achieves over 2 inference speedup with minimal degradation in output quality, offering an effective and generalizable solution for accelerating diffusion-based generative models.
@@ Introduction
Forward and reverse process in diffusion modes. Diffusion models gradually inject noise into data and learn to reverse this process in order to generate clean data from noise. In this work, we focus on the latent-space formulation of noise injection and denoising introduced by (Rombach et al. 2022), where the diffusion process is performed in the latent space. In the forward process, the posterior distribution of the noisy latent variable at timestep t admits a closed-form expression:
@@
When the number of selected blocks and the policy is applied within timesteps , Sortblock achieves a speedup on Flux.1-dev while maintaining comparable image quality in terms of both visual aesthetics and prompt alignment. Notably, different speed–quality trade-offs can be flexibly controlled by tuning and the application interval of the caching strategy.
@@
| Speedup | Latency (s) | VBench | SSIM | PSNR | LPIPs | |
| Wan2.1, 50 steps | 86 | 79.55% | - | - | - | |
@@
| Speedup | Latency (s) | VBench | SSIM | PSNR | LPIPs | |
| HunyuanVideo, 50 steps | 53 | 79.04% | - | - | - | |
@@ Discussion
Impact of reuse frequency. As illustrated in Fig. 6, we visualize how the generated images evolve with increasing values of . As the generation time decreases, the main semantic content of the image remains consistent, while the fidelity of fine-grained details gradually degrades. This observation is consistent with the insights presented in Fig. 1, which models the dynamic recomputation ratio over timesteps. Varying the value of provides a flexible control mechanism for balancing between generation speed and visual quality.

### v2

@@ Abstract
Diffusion Transformers (DiTs) have demonstrated remarkable generative capabilities, particularly benefiting from Transformer architectures that enhance visual and artistic fidelity. However, their inherently sequential denoising process results in high inference latency, limiting their deployment in real-time scenarios. Existing training-free acceleration approaches typically reuse intermediate features at fixed timesteps or layers, overlooking the evolving semantic focus across denoising stages and Transformer blocks.To address this, we propose SortBlock, a training-free inference acceleration framework that dynamically caches block-wise features based on their similarity across adjacent timesteps. By ranking the evolution of residuals, SortBlock adaptively determines a recomputation ratio, selectively skipping redundant computations while preserving generation quality. Furthermore, we incorporate a lightweight linear prediction mechanism to reduce accumulated errors in skipped blocks. Extensive experiments across various tasks and DiT architectures demonstrate that SortBlock achieves over 2 times inference speedup with minimal degradation in output quality, offering an effective and generalizable solution for accelerating diffusion-based generative models.
Project Page — https://sortblock.github.io
@@ Introduction
Forward and Reverse Process in Diffusion Models. Diffusion models gradually inject noise into data and learn to reverse this process in order to generate clean data from noise. In this work, we focus on the latent-space formulation of noise injection and denoising introduced by (Rombach et al. 2022), where the diffusion process is performed in the latent space. In the forward process, the posterior distribution of the noisy latent variable at timestep t admits a closed-form expression:
@@
When the policy refresh interval and the policy is applied within timesteps , SortBlock achieves a speedup on Flux.1-dev while maintaining comparable image quality in terms of both visual aesthetics and prompt alignment. Notably, different speed–quality trade-offs can be flexibly controlled by tuning and the application interval of the caching strategy.
@@
| Speedup | Lat.(s) | VBench | SSIM | PSNR | LPIP | |
| Wan2.1 (50 steps) | 86 | 79.55% | - | - | - | |
@@
| Speedup | Lat.(s) | VBench | SSIM | PSNR | LPIP | |
| HunyuanVideo (50 steps) | 53 | 79.04% | - | - | - | |
@@ Discussion
Impact of Reuse Frequency. As illustrated in Fig. 6, we visualize how the generated images evolve with increasing values of . As the generation time decreases, the main semantic content of the image remains consistent, while the fidelity of fine-grained details gradually degrades. This observation is consistent with the insights presented in Table. 1, which models the dynamic recomputation ratio over timesteps. Varying the value of provides a flexible control mechanism for balancing between generation speed and visual quality.
@@ Conclusion
Acknowledgements
This work is supported by the National Key Research and Development Program of China under Grant No. 2023YFB4604100.